EVA Report – April 4th

Crew 279 EVA Report 04-04-2023
EVA #02
Author: Antoine de Barquin Crew Commander
Purpose of EVA: Training EVA + communication test with crew members and with the hab.
Start time: 9:00
End time: 10:28
This EVA was also the first for the rest of the crew and was therefore also used as training. The EVA area was Marble Ritual as the day before. The objective was to learn how to correctly perform the EVA which includes several exercises such as the depressurization process, driving the rovers, radio communication and EVA end report. And on the other hand, to test different means of communication such as the designation of an object in the distance to show a direction and the language of the signs used in diving.
The crew scientist took advantage of the mobility exercises in wetsuit to deposit a dosimeter useful for his experience.
The crew finally decided to come back to the rover and start the return to the base at 10:10 am after having explored the area well. The reports were well done before entering the decompression chamber and the crew came back in time.
Destination: Marble Ritual
Coordinates (use UTM WGS 84): 519000 E ,4250500 N
Participants: 4: Antoine (Commander) (EVA Leader), Aglaé (Journalist), Ttele (HSO), Thomas (Scientist).
Road(s) and routes per MDRS Map: Cow Dung Road 0110
Mode of travel: First drive of Rover with spacesuits and accommodation with maneuver to the Marble Ritual then walk around no farther than 200 meters for communication check.

Operations Report – April 4th

Crew 279 Operations Report 04-04-2023
SOL: 2
Name of person filing report: Ioana Dimitrova
Non-nominal systems: robotic observatory and suit number 8
Notes on non-nominal systems: Robotic observatory: awaiting for a solution from the manufacturer
Suit number 8: Problem is being assessed by crew engineer (details in Summary of suits and radios)

Spirit rover used: YES
Hours: 213.9
Beginning charge: 100%
Ending charge: 87%
Currently charging: YES
Opportunity rover used: YES
Hours: 119.7
Beginning charge: 100%
Ending charge: 89%
Currently charging: YES
Curiosity rover used: no
Hours: 226.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Perseverance rover used: no
Hours: 260.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
General notes on rovers: none
Summary of Hab operations: Music influence on stress experiment was conducted. Teamwork exercise.
WATER USE: 52.28 (see note below)
Water (static tank): 432.37
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 14:00 to 21:00 for experiments and reports
Summary of suits and radios:
EVA crew reported a bit of fog forming in suit number 8 during EVA, not enough for the EVA to be dangerous or to be aborted. After testing, we noticed the suit ventilator’s airflow becomes less strong more rapidly than the others. The voltage of the battery also goes down quicker than the other two suits (suit type: helmet and suit combined type). The problem does not seem to be severe as the ventilator still works and clears the fog out but in a less optimal way than others. We suspect the battery is losing its condition. We will opt for other suits on the next EVA awaiting a repair on this one.
Radios charging nominal. Headset number 5 was replaced and labeled. It was tested and worked nominal, ready to use for the next EVA. Malfunctioning headset is in the RAM.
Summary of GreenHab operations:
Heater: Off
Supplemental light: Off
Harvest: none
Summary of ScienceDome operations: extremophiles were put in culture again because of contamination. EVA communication training. 3D printer reconstruction.
Dual split: Heat or AC, Off
Summary of RAM operations: no activity
Summary of any observatory issues: see above + cloudy weather and high winds so cupola was not opened
Summary of health and safety issues: none
Questions, concerns and requests to Mission Support:
A smell was detected on the lower deck. We located the smell coming from above the airlock that leads to the other parts of the station. We checked to see if the smell was coming from gas with the gas detector but the gas detector did not detect anything. We then tried touching and observing to see if there were any burnt cables or plastic caused eventually by a short circuit. After that, we called Mission Support who responded that this smell could occur, coming from the exhaust of the water heater when there are high winds.

GreenHab Report – April 4th

Crew 279 GreenHab Report 04-04-2023

GreenHab Officer: Augustin Tribolet
Environmental control: Open during day and heater during night
Average temperatures: 62.2 F
Hours of supplemental light: N/A
Daily water usage for crops: 12.6 gallons
Daily water usage for research and/or other purposes: N/A
Water in Blue Tank: 160.2 gallons
Time(s) of watering for crops: 06:15, 18:00
Changes to crops: N/A
Narrative: As part of the day-to-day activities of the GreenHab, general watering and maintenance of plants and crops was done, no new planting was performed. The door was left closed all day due to the strong wind and low sunlight.
Harvest: N/A
Support/supplies needed: N/A

Journalist Report – April 4th

CREW279 – Journalist Report 04APR2023

Journalist Report : Aglaé Sacré, crew journalist
SOL2: Surviving Mars temperature

Second day on the planet Mars!
This morning, we had a breakfast based on freeze-dried fruits! A real treat!
The other half of the crew went to discover the Martian lands in very cold temperatures, it is sure, we are on Mars.

After discovering the surroundings, Thomas, our scientist, started his experiment with Be.Sure, which consists in measuring the radiation doses received, so he buried his dosimeter which will serve as a reference. Each of us also wears a passive dosimeter to be able to process all the data at the end of the experiment.

Short lunch break, today Agnès and Ioana have been busy preparing bread for us! A real success with a superb soup!

After having eaten well, our commander and our astronomer gave us a course of communication to help with the next EVA, in a first time we learned to direct ourselves with simple signals by pointing at a point and by giving its distance. We also learned some simple signals from the diving techniques.

This afternoon the experiments are fully launched, what a joy to be able to work with such a view! Agnes has started to cultivate her extremophiles, they have already found their place on the planet Mars, they are ready to take up the challenge of surviving under the environmental conditions of the red planet. Thank you to BCCM for helping with this incredible biological experiment.

After dinner, our safety and health officer proposed a little team building at the beginning of the mission in order to spend a nice moment all together! A nice evening to end this SOL 2.

Sol Summary – April 4th

Crew 279 Sol Summary Report 04April2023

Sol: 2
Summary Title: Surviving Mars temperature
Author’s name: Agnès Dekeyser, Crew Executive Officer
Mission Status: Nominal
Sol Activity Summary:
After a good night of sleep, the crew was ready to start his second day on Mars planet. The mission began at 9 am with an EVA for the martionauts that didn’t get the training the day before. The purpose was again to practice maneuvering a rover and to check communications between each other and with the base. They also got training from their Commander to learn how to rapidly and precisely point a destination or an object. The same training was given to the rest of the Crew right after lunch, along with a training by the Crew Astronomer on how to communicate in the event of a communication technology malfunction.
For lunch, the crew enjoyed the bread baked that morning by the martionauts who had stayed in the Hab. The first bread of the mission. They have 10 sol left to improve their recipe.
The rest of the day was devoted to everyone’s experiment. Martionauts were testing the 3D printer, preparing map for radioactivity measure or for drone mapping, cultivating bacteria, observing the crew behavior, listening to music and analyzing the heartbeat, or answering anonymous questionnaire
Look Ahead Plan:
Tomorrow, an EVA is planned to begin the measure of radioactivity around the MDRS. The next EVA’s will be to progress on each other’s experiences. The crew will continue the radioactivity measures, perform several drone mapping, and expose bacteria to the MDRS environmental conditions.
Anomalies in work: The EVA suit number 8 needs to be repaired: poor ventilation, battery seems to lose its condition.
Weather: Cold and windy, high 6°C/43°F, low 0°C/32°F
Crew Physical Status: Nominal
EVA: Training
Reports to be filed: Journalist report, Greenhab report, Operations report, EVA report, EVA request.
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site